OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of 2/3?

OpenStudy (aroub):

y=mx+b ; m=2/3 , (6,3) 3=2/3(6)+b 3=4+b -1=b => y=2/3x-1

OpenStudy (anonymous):

But These Are The Solutions y =-1/2 x – 5 y =-1/2 x – 2 y =1/2 x + 2 y =1/2 x + 5

OpenStudy (aroub):

Okay, wait. Let me redo it :)

OpenStudy (aroub):

y-y1=m(x-x1) y-3=2/3(x-6) y-3=2/3x-4 y-2/3x=3-4 y-2/3x=-1 y=2/3x-1 Same answer..